Calculate the energy stored in 0.632 mH inductor with a current of 30.0 A.

  1. Calculate the energy stored in 0.632 mH inductor with a current of 30.0 A.
Answer #1

Magnetic energy stored in an inductor is U= (1/2)*L*I​​​​​​2

Where L is inductance of the inductor = 0.632 mH = 0.632×10-3 H.

And I is the current through the inductor = 30 A.

Energy stored U = (1/2)*(0.632×10-3)*(30)2 joule = 0.2844 joule.(ans)